OpenStudy (anonymous):

Solve 3/4(8p - 20) = 7

OpenStudy (debbieg):

is it \(\dfrac{3}{4}(8p - 20) = 7\) or \(\dfrac{3}{4(8p - 20)} = 7\)?

OpenStudy (anonymous):

the first one

OpenStudy (debbieg):

ok... so you need to solve for p, which means you need to isolate the p on the one side. The first thing I woulld do is get rid of that pesky fraction. What could you multiply by on both sides of the equation, to eliminate the fraction?

OpenStudy (anonymous):

i think the answer is 11/3 but im not sure

OpenStudy (anonymous):

4/3

OpenStudy (debbieg):

Yes, 4/3 will do it - or even just 4, then you don't get a fraction on the other side.

OpenStudy (debbieg):

Well, you can always check your solution in the original equation. If it makes it a true statement, then that's the solution. One of the beauties of solving equations. :)

OpenStudy (debbieg):

\[3(8p - 20) = 7\cdot4\]\[24p -60 = 28\]\[24p = 88\]\[p = \dfrac{88}{24}\]\[p = \dfrac{11}{3}\]

OpenStudy (debbieg):

Looks like the right answer to me. What about it were you unsure about?

OpenStudy (anonymous):

i just wanted to check my answer

OpenStudy (debbieg):

okeydoke. \(\Large \checkmark\) :)

OpenStudy (anonymous):

thanks!